REU 2006: SCHEDULE FOR WEEKS FIVE AND SIX

AFTERNOON TALKS ARE IN ECKHART 206 (unless posted otherwise) (May’s morning talks in R276)

Date  YSP  DeMarco  Babai  Souto  Unver  May
Mon, July 17 9:30 – 2:30 4:30 – 6:00  2:45 – 4:15
Tue, July 18 9:30 – 2:30  2:45 – 4:15 4:30 – 6:00
Wed, July 19 9:30 – 2:30 4:30 – 6:00  2:45 – 4:15
Thu, July 20 9:30 – 2:30  2:45 – 4:15 4:30 – 6:00
Fri, July 21 9:30 – 2:30 4:30 – 6:00  2:45 – 4:15
Mon, July 24 9:30 – 2:30 4:30 – 6:00 2:45 – 4:15  10:30 – 12:00
Tue, July 25 9:30 – 2:30  2:45 – 4:15 4:30 – 6:00
Wed, July 26 9:30 – 2:30 4:30 – 6:00 2:45 – 4:15  10:30 – 12:00
Thu, July 27 9:30 – 2:30  2:45 – 4:15 4:30 – 6:00
Fri, July 28 9:30 – 2:30 4:30 – 6:00 2:45 – 4:15  10:30 – 12:00

Here is the schedule in summary form:

FIFTH WEEK:

YSP 9:30 – 2:30 M-F DeMarco (Hausdorff dimension) 4:30 – 6:00 MWF Souto (Geometry) 2:45 – 4:15 TTh Unver (Algebraic geometry) 4:30 – 6:00 TTh May (K-theory) 2:45 – 4:15 MWF

SIXTH WEEK:

YSP 9:30 – 2:30 M–F DeMarco (Fractals and dimension) 4:30 – 6:00 MWF Babai (Discrete mathematics) 2:45 – 4:15 MWF Souto (Geometry) 2:45 – 4:15 TTh Unver (Algebraic geometry) 4:30 – 6:00 TTh May (K-theory for SESAME participants) 10:30 – 12:00 MWF

TUTORIALS and work sessions for SESAME participants during YSP work hours